OpenStudy (anonymous):

integral(-1/(x ln x))dx = Integral(1/y) dy

OpenStudy (anonymous):

integral(-1/(x ln x))dx = Integral(1/y) dy = ln(ln(x)) = ln(y) +c = ln(x) = y +c Is this correct?

hartnn (hartnn):

u missed a *minus* sign integral(-1/(x ln x))dx = Integral(1/y) dy =- ln(ln(x)) = ln(y) +c =- ln(x) = y +c

OpenStudy (anonymous):

oh... hehe apart from that sorry =)

hartnn (hartnn):

yes, apart from that its absolutely correct.

OpenStudy (anonymous):

because the site i am reading has "y = c ln x" And that doesn't seem right...?

hartnn (hartnn):

no, i don't think thats correct.

OpenStudy (anonymous):

He even misses the minus sign also. Thanks for confirming this for me.

hartnn (hartnn):

ok, welcome ^_^